OpenStudy (anonymous):

still having trouble on these and the steps , x+y=5 3x+2y=11

OpenStudy (unklerhaukus):

solve simultaneously?

OpenStudy (anonymous):

Do you know how to use substitution?

OpenStudy (anonymous):

x + y = 5 -y x = 5 - y Substitute the value of x into the second equation: 3x+2y=11 3(5 - y) + 2y = 11 15 - 3y + 2y = 11 combine like terms: 15 - y = 11 Solve for y: 15 - y = 11 -15 -y = -4 /-1 y = 4 x+y=5 x + 4 = 5 solve for x: x + 4 = 5 -4 x = 1 So x = 1 & y = -4 (1,-4)

OpenStudy (anonymous):

Hope this helps!

OpenStudy (anonymous):

omg yes it did helpp @jazy

OpenStudy (anonymous):

:)